### v1

# A Study of Plasticity Loss in On-Policy Deep Reinforcement Learning

Abstract
Continual learning with deep neural networks presents challenges distinct from both the fixed-dataset and convex continual learning regimes. One such challenge is plasticity loss, wherein a neural network trained in an online fashion displays a degraded ability to fit new tasks. This problem has been extensively studied in both supervised learning and off-policy reinforcement learning (RL), where a number of remedies have been proposed. Still, plasticity loss has received less attention in the on-policy deep RL setting. Here we perform an extensive set of experiments examining plasticity loss and a variety of mitigation methods in on-policy deep RL. We demonstrate that plasticity loss is pervasive under domain shift in this regime, and that a number of methods developed to resolve it in other settings fail, sometimes even resulting in performance that is worse than performing no intervention at all. In contrast, we find that a class of “regenerative” methods are able to consistently mitigate plasticity loss in a variety of contexts, including in gridworld tasks and more challenging environments like Montezuma’s Revenge and ProcGen.
1 Introduction
In many important machine learning domains, such as continual learning and online reinforcement learning (RL), training data are not wholly available simultaneously. Rather, in these settings, data arrive sequentially over a long period of time. Irrespective of the available quantity of training data, we would typically like to have the highest performing model possible at any given point. Unfortunately, these sequential learning scenarios are known to present optimization issues for neural networks. Specifically, if a model is naively updated to convergence each time new data arrive, and the amount of training data effectively increases, the resulting model generalizes worse than it would under more standard, non-sequential constraints. This phenomenon has been introduced by Ash and Adams, (2020) as the “warm-start problem,” where it was found that randomly initialized models generalize well but are expensive to fit, and warm-started models generalize poorly but converge far more quickly.
An adjacent problem to warm-starting is that of plasticity loss (Lyle et al.,, 2023), which presents itself both in the supervised learning and reinforcement learning settings. Like in the warm-start problem, plasticity loss describes a degradation in a model’s ability to fit new data as training progresses, ultimately harming both the sample efficiency and asymptotic performance. If an agent achieves worse performance when fitting first a source distribution and then a distinct target distribution than it would if instead trained on the target distribution in isolation, the degradation is attributed to plasticity loss.
Whereas the warm-start problem is restricted to test performance, plasticity loss is generally measured as a degradation in performance on training data, often under some form of environmental distribution shift. Despite the typical focus on training data, an ideal solution to plasticity loss would be able to prevent performance degradation both for data seen and unseen when fitting the policy. Importantly, the plasticity loss phenomenon is distinct from overfitting, as resolving plasticity loss should improve both training performance and generalization. In contrast, overfitting describes the opposite, where improving training performance has deleterious effects on generalization.
This work studies the plasticity loss phenomenon in detail for the on-policy reinforcement learning setting. We introduce three distinct kinds of distribution shift to facilitate our analysis as well as a variety of environments and tasks. While several interventions have already been proposed, we demonstrate that some of the methods which proved successful for addressing plasticity loss in the supervised learning or off-policy reinforcement learning setting fail in the on-policy setting or under our proposed forms of distribution shift. We further highlight several criteria that we believe are necessary to remedy the issue. Based on these insights, we describe several techniques that resolve plasticity loss on the environments we consider.
In summary, the contributions of this paper are:
  • We extend studies of plasticity loss and the warm-start problem to the on-policy regime, finding that they are a persistent issue across a variety of environmental distribution shift conditions.
  • We provide an in-depth analysis of the correlates of these pathologies, studying several types of environmental settings, model architectures, and previously proposed approaches for mitigation in settings related to on-policy reinforcement learning. Importantly, we include generalization trends in our consideration of these phenomena.
  • We characterize properties of methods that seem necessary for interventions to be successful at both addressing plasticity loss and ensuring generalization performance is maintained, and provide recommendations along these lines.
2 Preliminaries
An RL algorithm is described as “on policy” if it is trained using data collected from its own policy. While generally more sample inefficient than an off-policy alternative, where the data collection mechanism deviates from the policy that is being trained, on-policy methods have become more commonly used in practical RL problems because of their simplicity and reliability (Andrychowicz et al.,, 2020).
Proximal Policy Optimization (PPO) is a ubiquitous on-policy deep RL algorithm (Schulman et al.,, 2017). As a trust-region method, PPO constrains the extent by which a weight update can modify the current policy, quelling instability issues caused by high-variance policy gradients. Despite this robustness, on-policy algorithms still update on data corresponding only to the agent’s current experience, which may be suboptimal for capturing necessary information around solving the underlying MDP. There is preliminary evidence that plasticity loss can be present in on-policy RL learning problems (Dohare et al., 2023b, ). While prior work has demonstrated this effect in an ad hoc manner, we use multiple environments and distribution change conditions to systematically characterize plasticity loss—and how to resolve it—in the on-policy setting.
2.1 Simulating environmental distribution shift
This paper primarily studies three distinct forms of distribution shift that we can reasonably expect a plasticity-preserving intervention to mitigate. Each experiment is organized into $r$ distinct rounds. In each round we supply $k$ environments from some distribution to the agent, fit a policy for that set of environments, and measure performance at the end of the round. To induce distribution shift, between each round we modify the data according to one of the three strategies described below. The environmental modifications we consider are:
Permute. Input pixels are randomly shuffled from round to round. Each of the $k$ environments in a given round are shuffled in the same way, but distinct from the shuffling in any other round. This method is used in Dohare et al., 2023b . We expect that an optimal learning algorithm will be able to achieve equivalent performance on the initial and subsequent rounds.
Window. We supply the agent with $k$ new environments for the same task. Here previously seen environments are no longer accessible, so the agent can only optimize with respect to the current $k$. This approach is considered by Abbas et al., (2023) and Lyle et al., (2023). We expect that an optimal learning algorithm will be able to achieve equivalent or better performance on subsequent rounds as compared to the initial round.
Expand. Like in the window modification, $k$ new environments are supplied to the agent. Here, however, they are appended to the total amount of training data, such that the number of available environments for training at round $n$ is $k\times n$. This is most similar to the warm-start environment studied primarily in Ash and Adams, (2020) and revisited in Igl et al., (2020). In this setting we expect that even for an optimal learning algorithm training performance to decrease each round, as the dataset becomes more difficult to fit. Correspondingly, we expect the generalization performance to increase as the training distribution expands to better capture the true distribution.
In the case of the latter two modifications, data at each new round are drawn from the exact same distribution as environments from previous rounds. As we demonstrate in Section 4.1, these still induce significant performance degradation in warm-started models. This observation is surprising—we often think of a model’s initialization as a sort of “prior” over learned functions, and here we demonstrate that even an initialization obtained by training on data distributed identically to the current round still hinders performance in on-policy learning.
We use a simple gridworld task as our main sandbox for studying plasticity loss. The environment is drawn from the NeuroNav library (Juliani et al.,, 2022), and the goal of the agent is to collect rewarding blue jewels while avoiding punishing red jewels in a fixed time window (100 time-steps per episode). Each sample of the environment from the distribution of possible tasks changes both the location of the jewels and the walls of the maze. Figure 1 shows environment modification examples.
3 Existing Approaches
A number of methods have been proposed to address plasticity loss in deep neural networks, though none were designed with on-policy reinforcement learning explicitly in mind. Instead, these methods have been primarily validated in the context of either supervised learning or off-policy RL. Each of these techniques can be divided into one of two groups: Interventions which are performed intermittently during training, and those which are applied continuously, either as part of the model architecture or as part of the model update process. For additional implementation details see Appendix Section A.
3.1 Intermittent Interventions
Figure 1: Examples of gridworld environment tasks. The agent (black triangle) begins each episode in the center of the environment. Blue jewels provide +1 reward, red jewels provide -1 reward, and dark grey walls prevent movement. Objects are placed randomly.
We consider an intervention to be intermittent if it is applied only at specific points during training. With one exception, each time the training data change, the interventions in this category are applied and training proceeds. Note that this implies that there must be awareness of when the training distribution shifts, which in practice may be difficult or infeasible.
Resetting final layer. Proposed by Nikishin et al.,, this method involves periodically replacing the weights of the final layer in the network with newly initialized values. It was demonstrated that this alleviates plasticity loss in certain off-policy RL.
Shrink+Perturb. This technique periodically reduces the magnitude of all weights in the network by a factor and then adding noise sampled from a new weight initialization which is scaled down by a separate factor (Ash and Adams,, 2020). It was demonstrated that this improves performance in batched continual learning settings. In our implementation the weight of these two factors are entangled such that they sum to one.
Plasticity Injection. Plasticity Injection replaces the final layer of a network with a new function that is a sum of the final layer’s output and the output of a newly initialized layer subtracted by itself (Nikishin et al.,, 2023). The gradient is then blocked in both the original layer and the subtracted new layer. It was shown that this method increases performance of off-policy RL agents in the Atari Learning Environment (ALE).
ReDo. This technique resets individual neurons withing the network based on a dormancy criteria at fixed intervals (Sokar et al.,, 2023), leading to improvement in performance in the context of off-policy RL agents trained on the ALE. Following Sokar et al.,, we reset the model parameters more frequently than we apply environmental distribution shifts.
3.2 Continuous Interventions
We characterize a method as continuous if it is applied at every step of optimization. Continuous interventions are desirable because they do not require an awareness of when a distribution shift has occurred, which can be challenging in practical scenarios where we want to mitigate plasticity loss.
L2 Norm. Discussed by Lyle et al.,, this method involves regularizing the network using the L2 norm. It was demonstrated that in certain continual learning settings this reduces plasticity loss.
LayerNorm. LayerNorm (Ba et al.,, 2016) is a ubiquitous deep learning regularization technique, and was shown useful for plasticity loss by Lyle et al.,. The authors demonstrated a performance improvement in the ALE when using off-policy RL.
CReLU activations. Proposed by Abbas et al., and studied further by Lee et al.,, this method involves replacing ReLU activations with the CReLU activation function (a concatenation of $\mathrm{ReLU}(x)$ and $\mathrm{ReLU}(-x)$ for an input scalar $x$). This ensures that the gradient is non-zero for all units in a given layer, which appears to mitigate plasticity loss when periodically switching between tasks in the ALE with off-policy methods.
Regenerative Regularization. Regenerative regularization uses an L2 penalty in the optimized loss to encourage model weights to be near their initial values (Kumar et al.,, 2023). It has been shown to be effective in a number of continual learning settings.
4 Experiments
Figure 2: Performance in the gridworld environment under each of the three distribution shift conditions. The degradation between rounds is evidence of plasticity loss. (A): Epoch-level training performance for permute modification. Dotted vertical lines indicate the end of each round, before a new environmental distribution shift is applied. (B): Round-level training performance for the Permute modification. Data points correspond to normalized mean reward in final 50 episodes of round. (C, D): Round-level training performance for the Window and Expand conditions. Top row: Training performance. Bottom row: Test performance.
Below we present results of a suite of experiments conducted using a 2D gridworld environment (Juliani et al.,, 2022), the procedurally generated CoinRun environment (Cobbe et al.,, 2019), and the Atari game Montezuma’s Revenge (Bellemare et al.,, 2013). All models are trained using PPO (Schulman et al.,, 2017). For the gridworld experiments we use an MLP encoder, and for the CoinRun and Montezuma’s Revenge experiments we instead use a convolutional encoder. We set the number of environment instances ($k$) to 100 for all experiments conducted with the gridworld and CoinRun environments. Like in previous sections, many of these results show reward trends as a function of the number of rounds, with each round denoting a new shift in the state distribution. These curves are shown in terms of normalized reward (i.e. we subtract the reward obtained after the first round from the reward obtained after each subsequent round) to compensate for the fact that several algorithms have a regularization effect that improves performance on average but does not necessarily improve the performance trend—it is possible to improve generalization, for example, without remedying the degradation in train or test performance induced by plasticity loss. Training details can be found in Appendix Section B.
4.1 Plasticity loss is apparent in on-policy RL
We find evidence of plasticity loss in the on-policy reinforcement learning setting under multiple kinds of distribution shift. Figure 2 presents these results. As mentioned earlier, for round-level plots like these, reward values are normalized such that the mean reward at the end of the first round is set to zero. As such, positive and negative values correspond respectively to increases and decreases in performance relative to the first round. Decreasing normalized reward indicates a decay in relative performance and is indicative of plasticity loss.
For both the permute and window shift conditions, Figure 2 demonstrates clear degradation in mean episodic reward for the warm-started model as compared to a model which has all of its weights reset between each round. While there is some positive transfer for the warm-start method in the expand shift condition, we find generalization performance suffers here in comparison to reset-all. We provide statistical tests comparing these methods to warm-start and random initialization baselines in Appendix Section D.
4.2 Predictors of plasticity loss and generalization
Figure 3: Top: Correlation plots of normalized mean reward in the gridworld environment compared against various model attributes. Each point is averaged over five replicates, and shows the final values produced after a ten-round experiment. Values for each measurement are normalized by its baseline level at initialization. Measurements that significantly correlate ($p<0.05$) with normalized reward are bolded. First row: Training performance. Second row: Test performance. Bottom: Values of the three predictive attributes during the course of training for each intervention and window change condition as compared to training performance. Final values of plots like these correspond to a single point in the correlation plots in the first two rows.
Identifying the exact statistical underpinnings of plasticity loss is an active area of research in continual and reinforcement learning. Here we identify several factors which are significantly correlated with plasticity loss as well as a degradation in generalization performance. Figure 3 presents correlation plots of normalized reward with various attributes of interest.
Overall, we find that weight magnitude and number of dead units are both significantly correlated with the normalized reward [$p<0.05$]. Here, “dead units” refers to ReLU activations that never produce non-negative outputs. Note that the LayerNorm and CReLU methods are not included in this analysis, due to their immediate impact on the weight magnitude and dead unit counts, respectively. For plots of these measurements over time during training, see Supplemental Figure 9.
We find that neither the gradient magnitude nor the magnitude of weight difference between updates has a significant relationship with either plasticity or generalization. This is in contrast to the findings of Abbas et al., and Nikishin et al.,, each of which examined plasticity loss in the off-policy rather than on-policy setting. The former found that weight difference between SGD updates is correlated with loss of plasticity while the latter found that gradient norms were predictive.
We also fit a generalized linear model (GLM) using all five criteria as independent variables and normalized reward as the dependent variable. Here we find that both gradient and weight magnitudes are significantly predictive of train performance [$p<0.01$], while dead unit count no longer is [$p>0.05$]. This suggests that the predictive power of the dead unit count can be accounted for by the weight magnitude. An additional GLM was fit to the test data, and we find that only the dead unit count remains predictive [$p<0.05$].
Although these analyses do not establish causality, the robust significance of weight magnitude in predicting plasticity loss suggests that plasticity loss may be a function of how much trainable parameters deviate from their initialization distribution. As such, successful mitigation strategies might curb unbounded increases in the weight norm of network parameters over the course of learning. This has the downstream effect of minimizing the number of dead units in the network. Below we compare mitigation strategies, finding that those that keep weight norm from excessively growing seem to perform best.
4.3 Architectural interventions do not fully address plasticity loss
Figure 4: Performance of intervention methods compared to baselines of warm-start and reset-all on the gridworld environment. Mean reward is normalized by performance at end of first round. Top row: Training performance. Bottom row: Test performance.
Recently a number of strategies have been introduced to specifically address plasticity loss via changes to the network architecture. These methods were previously validated in the off-policy setting only. Here we consider the behavior of two of these methods as it applies to the on-policy setting: CReLU (Abbas et al.,, 2023) and Plasticity injection (Nikishin et al.,, 2023). We find that both methods fail to address plasticity loss in the permute shift condition. The CReLU method is able to address plasticity loss in the window and expand settings only, while the plasticity inject method underperforms even the warm-start baseline in all three contexts (Figure 4).
4.4 Regularization methods do address plasticity loss
We next consider the class of methods which regularize the weights of the network, either continuously or intermittently. These methods include final layer reset (Nikishin et al.,, 2022), shrink+perturb (Ash and Adams,, 2020), a continuous “soft” variant of shrink+perturb (Dohare et al., 2023b, ), L2 regularization, ReDo (Sokar et al.,, 2023), and regenerative regularization (Kumar et al.,, 2023). We find that the final-layer reset fails to address plasticity loss in any of the three conditions. This is likely because the learned penultimate layer representation is insufficient for mapping states to optimal actions after a shift in state distribution. In contrast, the other five methods mitigate both plasticity loss and the warm-start problem to some extent in all three shift conditions (Figure 4). Of these methods, ReDo performs the worst as it is unable to benefit from the positive transfer which is possible in the cases of the “window” and “expand” conditions, as methods that retain information from previous rounds effectively have access to more training data in these conditions. ReDo underperforms even the warm-start baseline in the “expand” condition.
4.5 LayerNorm addresses training plasticity loss but not generalization
We also consider the effect of LayerNorm on plasticity loss. It was previously demonstrated in the off-policy setting that LayerNorm is able to mitigate some effects of plasticity loss (Lyle et al.,, 2023). Given that LayerNorm tends to improve learning across a number of contexts (Ba et al.,, 2016), it is important to attempt to separate the effects of LayerNorm on performance generally from its effect on plasticity specifically. We find that LayerNorm resolves plasticity loss in terms of training performance, but is inconsistent in its effect on generalization performance. For example, the LayerNorm model significantly underperforms even the warm-start baseline on the test data of the permute shift condition. We find however that combining soft shrink+perturb or regenerative regularization with LayerNorm results in overcoming both training plasticity loss and deleterious generalization trends (Figure 4).
4.6 Plasticity loss in CoinRun
We also use the CoinRun environment from the ProcGen suite of tasks to evaluate the set of candidate interventions (Cobbe et al.,, 2019, 2020). ProcGen is built using procedural generation (Cobbe et al.,, 2020), making it possible to study the same three distribution-shift conditions which were considered in the gridworld experiments. Here we present performance results for each of the intervention methods on the CoinRun environment.
Figure 5: Performance of intervention methods compared to baselines of warm-start and reset-all on the CoinRun environment. Mean reward is normalized by performance at end of first round. Top row: Training performance. Bottom row: Test performance.
Overall, we find results largely consistent with earlier results—there is significant plasticity loss present in all three tasks as measured by the gap between the warm-start and reset-all conditions. We also find that reset-final, CReLU, and plasticity injection are all unable to fully resolve the pathology. In contrast, the three most effective methods in the gridworld task (LayerNorm, regenerative regularization, and soft shrink+perturb) are all effective in CoinRun as well. See Figure 5 for plots of the normalized mean rewards over the course of training for all methods. For results on two additional ProcGen environments, see Appendix D.
For round-level performance plots of all methods, see Section D in the appendix. We find that relative efficacy of the intervention methods is consistent across ProcGen environments.
4.7 Plasticity loss in Montezuma’s Revenge
Figure 6: A comparison between RND agents trained with and without a plasticity-loss mitigating intervention. Experiments are done over twenty replicates and shaded regions show standard error.
Montezuma’s Revenge is an Atari game that is often used to benchmark the quality of exploration procedures in RL (Bellemare et al.,, 2013; Salimans and Chen,, 2018). The environment consists of many separate rooms, some only accessible via a locked door. The agent must avoid obstacles, find and use keys, and traverse several rooms before obtaining non-trivial reward. Because of this reward sparsity, exploration is needed to incentivize the agent to advance beyond the periphery of its experience.
Montezuma’s Revenge is designed such that the agent only sees the room it currently occupies. Accordingly, the distribution over input states expands as the agent learns to explore the environment. Unlike other ALE games, this introduction of significantly different states as training progresses induces a level of distribution shift. Within our framework, the distribution shift of Montezuma’s Revenge is most similar to the “expand” condition studied earlier.
In this experiment we consider two versions of a policy trained using the Random Network Distillation (RND) exploration bonus (Burda et al.,, 2018). All architectures and hyperparameters are inherited from the original RND paper. Figure 6 compares a typical RND policy to ones that are trained in conjunction with the soft shrink+perturb or regenerative regularization interventions. The plasticity loss mitigating policy achieves higher reward than a corresponding agent fitted in a standard fashion. Due to computation constraints, we did not evaluate the other intervention methods.
5 Discussion
Plasticity loss has been identified and rigorously studied in the continual learning and off-policy reinforcement learning settings. We find that it is also an issue in the on-policy setting under a variety of different conditions. Methods introduced in the off-policy setting such as plasticity injection and CReLU appear to be not as consistently effective in the on-policy setting. We hypothesize that this degradation may be due to the violation of the iid assumption which supervised learning and off-policy reinforcement learning rely on. Further, these methods require architectural modifications of the policy and value networks, and are thus less general than regularization-based alternatives.
In contrast, the class of regularization methods, which include L2 regularization, ReDo, shrink+perturb, and regenerative regularization all are able to significantly mitigate plasticity loss in our experiments. Of these methods, soft shrink+perturb displayed the best generalization performance, and can be easily combined with layer normalization. LayerNorm, a now ubiquitous method in supervised learning, was previously introduced to address plasticity loss in the off-policy setting (Lyle et al.,, 2023), and we find that it is also effective in the on-policy setting. Due to its general regularizing effects, LayerNorm increases performance in the absence of distributional shift (Ba et al.,, 2016), suggesting that it is a generally useful method for on-policy reinforcement learning methods such as PPO.
A consistent feature of the class of regularization methods discussed here is that they all normalize the network parameters towards their initial distribution. This has the effect of decreasing the weight magnitude and reducing the number of dead or saturated activation units, as seen in Figure 3. It has recently been proposed that both of these measurements may be a proxy for the underlying curvature of the optimization landscape (Lewandowski et al.,, 2023); studying this connection could be a fruitful avenue for future research.
6 Related Work
In the context of generalization in continual learning settings, a related issue to plasticity loss is the warm-start problem (Ash and Adams,, 2020). Recent work has suggested that the benefits of methods like shrink+perturb to generalization may depend on the presence of noise in the labels used for training (Zaidi et al.,, 2023). In the absence of label noise, other regularization techniques were found to be competitive with methods that make use of re-initialization.
Another perspective on the problem of loss of plasticity has focused on the content of early learning in a network.  Achille et al., (2018) demonstrate that neural networks have critical early periods of sensitivity to training data which are analogous to those in animals (Hensch,, 2004). In the biological context this sensitivity is tied directly to neuronal plasticity. This connection to critical period learning was also drawn in work on off-policy RL, in which a “primacy bias” was proposed to explain the loss of plasticity seen in a number of task domains (Nikishin et al.,, 2022). Utilizing this insight, the periodic soft-resetting of the neural network has enabled significant increases in the sample efficiency of off-policy RL methods (D’Oro et al.,, 2022; Schwarzer et al.,, 2023).
In the deep RL domain a similar problem called ray interference has been identified (Schaul et al.,, 2019). Ray interference arises from the conflicting gradient signals related to unique sub-tasks required to solve a single more complex task in online deep reinforcement learning. It is possible that for complex tasks such as CoinRun or Montezuma’s Revenge there are unique sub-tasks which are learned and whose gradients may interfere with one another. Exploring the distinction between ray interference and plasticity loss in the RL setting is a promising future direction.
In the off-policy RL setting the problem of plasticity loss has also been studied under the name of capacity loss (Lyle et al.,, 2022). While plasticity loss refers to a property of the network-task interaction, capacity loss is a more generic term to refer to the properties of a network that are task invariant. This work advocates for a solution similar to that of Kumar et al., (2023), and can thus be considered in the class of methods which regularize the network towards an initialization distribution.
Within the domain of on-policy RL there is a related phenomena to plasticity loss called policy collapse, which refers to a degradation of performance as training progresses, even on a fixed dataset (Dohare et al., 2023c, ). Although these are distinct phenomena, there is likely overlap in the success of mitigation methods. In terms of direct studies of plasticity loss in on-policy RL, both Dohare et al., 2023a and Igl et al., (2020) demonstrate plasticity loss in the on-policy setting, and propose a method to address it.
As part of their analysis, Dohare et al., 2023b propose a plasticity-enhancing method called Continual Backprop, which is another form of regularization towards the initialization distribution, however it is one which is selective on a per-neuron basis. The approach is similar to ReDo (Sokar et al.,, 2023), which we analyze here. It is also related to DrM (Xu et al.,, 2023), which builds on the DrQ algorithm.
It is finally worth mentioning that notions of plasticity loss have been studied in the neuroscience community as well. These range from classic work demonstrating the inability of animal models to fully accommodate new sensory input after developing cognitively without it (Wiesel and Hubel,, 1963), to more recent work characterizing plasticity loss in humans as a causal factor in various psychopathologies (Peled,, 2005; Carhart-Harris et al.,, 2023; Juliani et al.,, 2024).
7 Conclusion
This work studies the problem of plasticity loss in the context of on-policy deep reinforcement learning. Like in continual learning and off-policy RL, we find plasticity loss is a pervasive issue across a number of different environments and forms of distribution shift. We further find that some methods previously proposed to resolve the issue in other problem settings, such as CReLU and plasticity injection, do not transfer to the on-policy regime. The methods which best resolve plasticity loss are those which act as continual regularizers as opposed to intermittent interventions. Within this class of methods, soft shrink+perturb and regenerative regularization seem to perform best, particularly when used in conjunction with layer normalization.
Appendix A Implementation Details
Below are the specific implementation details for each intervention. Wherever possible we followed the methods provided by the papers which introduced each method. Relevant hyperparameters used are listed in Table 1.
Shrink+perturb. When the intervention is applied all learnable parameters in the network are iterated through and scaled by $\alpha$. All parameters are then additively combined with newly sampled initialization parameters which are scaled by $\beta$. For a set of parameters $x$, this corresponds to $x_{new}=\alpha x_{current}+\beta x_{init}$, where $x_{init}$ is sampled from the parameter initialization distribution. For all experiments $\alpha=1-\beta$.
Soft shrink+perturb. The shrink+perturb procedure is applied after each step of gradient descent instead of only at specific intervals. Specifically, for a set of parameters $x$, this corresponds to $x_{new}=\alpha x_{current}+\beta x_{init}$, where $x_{init}$ is sampled from the parameter initialization distribution. For all experiments $\alpha=1-\beta$.
CReLU. All ReLU activations in the network are replaced with CReLU. The equation for CReLU is: $y=concat(ReLU(x),ReLU(-x))$. We use the method where input parameter sizes are halved compared to the baseline so that output parameters remain consistent for each layer. This results in a network with half the number of parameters for the effected layers.
Regenerative Regularization (L2). An additional loss term is added to the PPO update in each step of gradient descent. This term corresponds to L2 norm of the difference between the current parameters of the network and the network parameters at initialization. This loss is scaled by $\alpha$.
Plasticity injection. We follow the method described by Nikishin et al., for the initial intervention, only performing plasticity injection on the final layers of the network (i.e. value and policy heads). To ensure computational efficiency, during subsequent interventions we combine the previous parameter values into a single set of weights and biases (i.e. $w_{old}=w_{old}+w_{newa.}w_{newb}$). This ensures that at any given time the plasticity-injected layer is represented by $y=x_{old}+x_{newa.}sg(x_{newb})$.
LayerNorm. We apply the LayerNorm Ba et al., (2016) function before the ReLU activation at each layer.
Reset-final. At each intervention point we replace the weights of the final layers (i.e. value and policy heads) of the network with freshly initialized values.
L2 Norm. We apply an additional loss term to the PPO update which corresponds to the L2 norm of the weights of the network. This loss is scaled by a parameter $\alpha$.
ReDo. At a fixed epoch interval the procedure from Sokar et al., is performed on all layers of the neural network. A mini-batch of training data is used to calculate the dormancy rate for each unit, and units which are below $\tau$ are re-initialized.
Appendix B Environment and Model Details
Additional hyperparameters used for Gridworld and CoinRun models are provided in Table 2. Unless indicated otherwise, all experiments involve running five replicates with unique random seeds each. All experiments are conducted using either a single P100 or V100 GPU on a cloud machine.
Gridworld. The gridworld environment consisted of an 11x11x4 one-hot encoded observation space. Each observation was flattened into a single vector and provided as input to the model. In the permute environment variation, observations were split into 1x1x4 patches and randomized according to a fixed map within each round. The PPO model consisted of an MLP with two hidden layers, ReLU activations, and a “dual-head” architecture, as described in Schulman et al., (2017). The action space consisted of four possible actions, corresponding to movement in the cardinal directions. All episodes lasted 100 time-steps. Each round of training consists of 20,000 iterations of inference and model updates. Given ten rounds, the total number of iterations is 200,000.
CoinRun. The CoinRun environment consisted of a 64x64x3 observation space. This was provided to the PPO model as a tensor, and processed by a convolutional encoder with four layers and ReLU activations. Each convolutional layer utilized a kernel size of four and a stride of two. In the permute environment variation, images were broken into 8x8x3 patches and the position of patches was randomized according to a fixed map within each round. A “dual-head” architecture was also used. The action space consisted of the standard nine possible actions provided by the ProcGen environment Cobbe et al., (2020). Episode length varied based on agent behavior. Each round of training consists of 5,000 iterations of inference and model updates. Given ten rounds, the total number of iterations is 50,000.
Montezuma’s Revenge. Like in CoinRun, Montezuma’s revenge states are processed into observations of size 64x64x3 and passed through a convolutional encoder. We inherit hyperparameters from Burda et al., (2018), which were previously optimized for final performance. These include 4 randomly masked update epochs per observation, an entropy coefficient of 0.001, $\gamma_{E}$ = 0.999 and $\gamma_{I}$ =0.99. We use a learning rate of 0.0001 and 128 simultaneous agents simultaneously. All model updates are performed via the Adam optimizer. We use a shrinkage coefficient $1-10^{-8}$ and a perturbation coefficient of $10^{-8}$.
Appendix C Hyperparameters
Method Gridworld CoinRun
L2-Norm ($\alpha$) 1e-3 1e-3
Regen-Reg ($\alpha$) 1e-4 1e-2
Soft-S+P ($\beta$) 1e-6 1e-6
S+P ($\beta$) 0.5 0.5
ReDo period 10 epochs 10 epochs
ReDo ($\tau$) 0.025 0.025
Table 1: Optimal hyperparameter values used for different environments and interventions. Chosen values are the result of performing a sweep over possible values using permute environment shift condition.
Parameter Value
Number of hidden units 256
Learning rate 5e-4
Default activation ReLU
Minibatch size 64
Buffer size 1024
Discount ($\gamma$) 0.99
GAE parameter ($\lambda$) 0.95
Clip parameter ($\epsilon$) 0.2
Entropy parameter ($\beta$) 0.02
Number of update epochs 3
Table 2: Default values used in PPO algorithm for all gridworld and CoinRun experiments.
Appendix D Extended Results
Figure 7: Performance of intervention methods compared to baseline (warm-start) on Jumper task. Top row: Training distribution performance. Bottom row: Test distribution performance.
Figure 8: Performance of intervention methods compared to baseline (warm-start) on Fruitbot task. Top row: Training distribution performance. Bottom row: Test distribution performance.
Figure 9: Effect of learning under various conditions on five diagnostic metrics of interest.
Dep. Variable: train_reward No. Observations: 27
Model: GLM Df Residuals: 21
Model Family: Gaussian Df Model: 5
Link Function: Identity Scale: 0.13264
Method: IRLS Log-Likelihood: -7.6473
Deviance: 2.7855
Pearson chi2: 2.79
No. Iterations: 3 Pseudo R-squ. (CS): 0.5624
coef std err z P>|>|> |z|||| [0.025 0.975]
const -0.0920 0.118 -0.783 0.434 -0.322 0.138
entropy 1.6302 0.940 1.734 0.083 -0.213 3.473
weight_mag -1.8262 0.637 -2.868 0.004 -3.074 -0.578
weight_diff -35.4734 133.365 -0.266 0.790 -296.865 225.918
grad_norm 0.5742 0.184 3.116 0.002 0.213 0.935
dead_units 0.5074 5.905 0.086 0.932 -11.066 12.081
Table 3: Results of GLM using various diagnostic metrics to predict reward during training.
Dep. Variable: test_reward No. Observations: 27
Model: GLM Df Residuals: 21
Model Family: Gaussian Df Model: 5
Link Function: Identity Scale: 0.19798
Method: IRLS Log-Likelihood: -13.054
Deviance: 4.1576
Pearson chi2: 4.16
No. Iterations: 3 Pseudo R-squ. (CS): 0.5215
coef std err z P>|>|> |z|||| [0.025 0.975]
const 0.3249 0.144 2.263 0.024 0.043 0.606
entropy -1.4060 1.149 -1.224 0.221 -3.657 0.845
weight_mag 1.1259 0.778 1.447 0.148 -0.399 2.651
weight_diff 327.8923 162.934 2.012 0.044 8.547 647.238
grad_norm -0.0568 0.225 -0.252 0.801 -0.498 0.384
dead_units -21.6282 7.214 -2.998 0.003 -35.768 -7.488
Table 4: Results of GLM using various diagnostic metrics to predict reward on the test distribution.
Gridworld (Permute - Train)
Method t-value p-value
Warm-start t(7) = 2.524 0.04
CReLU t(7) = 10.477 0.0
Injection t(8) = 6.21 0.0
Reset-final t(7) = 2.444 0.044
L2 norm t(7) = -0.08 0.939
Shrink+perturb t(8) = -0.087 0.933
Soft shrink+perturb t(8) = 0.472 0.65
Regen reg (L2) t(8) = 0.134 0.897
LayerNorm (LN) t(8) = 0.787 0.454
Soft shrink+perturb + LN t(8) = -0.212 0.837
Regen reg (L2) + LN t(8) = -1.368 0.208
Table 5: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (permute) condition. T-values below zero correspond to train distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Gridworld (Window - Train)
Method t-value p-value
Warm-start t(8) = 0.759 0.469
CReLU t(7) = -1.563 0.162
Injection t(8) = 1.741 0.12
Reset-final t(8) = -0.017 0.987
L2 norm t(8) = -1.291 0.233
Shrink+perturb t(8) = -1.361 0.211
Soft shrink+perturb t(7) = -0.222 0.831
Regen reg (L2) t(8) = -1.068 0.317
LayerNorm (LN) t(7) = -2.107 0.073
Soft shrink+perturb + LN t(8) = -1.549 0.16
Regen reg (L2) + LN t(8) = -3.106 0.015
Table 6: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (window) condition. T-values below zero correspond to train distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Gridworld (Expand - Train)
Method t-value p-value
Warm-start t(8) = -0.926 0.382
CReLU t(8) = -2.247 0.055
Injection t(8) = -0.715 0.495
Reset-final t(8) = -1.301 0.229
L2 norm t(8) = -2.252 0.054
Shrink+perturb t(8) = -2.476 0.038
Soft shrink+perturb t(7) = -1.788 0.117
Regen reg (L2) t(8) = -2.879 0.021
LayerNorm (LN) t(8) = -4.098 0.003
Soft shrink+perturb + LN t(8) = -3.683 0.006
Regen reg (L2) + LN t(8) = -5.44 0.001
Table 7: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (expand) condition. T-values below zero correspond to train distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Gridworld (Permute - Test)
Method t-value p-value
Warm-start t(7) = 3.0 0.02
CReLU t(7) = 6.975 0.0
Injection t(8) = 3.595 0.007
Reset-final t(7) = 2.445 0.044
L2 norm t(7) = 3.009 0.02
Shrink+perturb t(8) = 0.907 0.391
Soft shrink+perturb t(8) = 2.177 0.061
Regen reg (L2) t(8) = 0.007 0.995
LayerNorm (LN) t(8) = 6.959 0.0
Soft shrink+perturb + LN t(8) = 0.235 0.82
Regen reg (L2) + LN t(8) = 0.804 0.445
Table 8: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (permute) condition. T-values below zero correspond to test distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Gridworld (Window - Test)
Method t-value p-value
Warm-start t(8) = 0.882 0.404
CReLU t(7) = 0.239 0.818
Injection t(8) = 1.444 0.187
Reset-final t(8) = 0.804 0.445
L2 norm t(8) = 0.216 0.834
Shrink+perturb t(8) = -0.781 0.458
Soft shrink+perturb t(7) = -0.046 0.965
Regen reg (L2) t(8) = -0.187 0.856
LayerNorm (LN) t(7) = -0.813 0.443
Soft shrink+perturb + LN t(8) = -2.349 0.047
Regen reg (L2) + LN t(8) = -1.714 0.125
Table 9: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (window) condition. T-values below zero correspond to test distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Gridworld (Expand - Test)
Method t-value p-value
Warm-start t(8) = 1.227 0.255
CReLU t(8) = 0.891 0.399
Injection t(8) = 1.815 0.107
Reset-final t(8) = 1.457 0.183
L2 norm t(8) = -0.512 0.623
Shrink+perturb t(8) = -1.021 0.337
Soft shrink+perturb t(7) = -0.527 0.614
Regen reg (L2) t(8) = 0.333 0.748
LayerNorm (LN) t(8) = 0.438 0.673
Soft shrink+perturb + LN t(8) = -1.793 0.111
Regen reg (L2) + LN t(8) = -0.579 0.579
Table 10: Results of statistical tests comparing different methods to baseline of reset-all in gridworld (expand) condition. T-values below zero correspond to test distribution performance better than baseline. T-values above zero correspond to performance worse than baseline. P-values above 0.05 correspond to methods which produce mean normalized rewards not significantly different from the baseline.
Figure 10: Performance of intervention methods compared to baseline (warm-start) on Gridworld task. Each error bar corresponds to normalized mean episodic reward. Tow row: Training distribution performance. Bottom row: Test distribution performance.
Figure 11: Performance of intervention methods compared to baseline (warm-start) on CoinRun task. Each error bar corresponds to normalized mean episodic reward. Top row: Training distribution performance. Bottom row: Test distribution performance.